The word problems is as follows:
The owners of a candy store wants to sell, for $5 a pound, a mixture of chocolate covered raisins, which usually sells for $3 per pound, and chocolate covered nuts, which usually sells for $8 per pound. They have a 50 pound barrel of the raisins. How many pounds of the nuts should they mix with the barrel of raisins so that they hit their target value of $5 per pound for the mixture?
Please help me I just think that I'm over thinking it.

Answer by saifoo.khan(162)

Hello,,
Let x = chocolate-covered raisins and y = chocolate covered macadamia nuts
5(x + y) = 3x + 8y
We are given that they have 50 pounds of chocolate covered raisins
5(50 + y) = 3(50) + 8y
250 + 6y = 150 + 8y
100 = 2y
y = 50 pounds

Thanks! ^_^
--------------------------------
einstein.saifoo@gmail.com